# Fairness-in-the-Workflow: How Machine Learning Practitioners at Big Tech Companies Approach Fairness in Recommender Systems

### v2

@@ 3. Methods
To understand how RS practitioners approach fairness in practice, we conducted 11 semi-structured interviews over a three-week period in 2022. We interviewed RS practitioners from a variety of large companies that deploy highly visible RS.
@@
We recruited participants via social media, direct emails to individuals in our networks or who we had identified based on publicly available information as working in RS, and snowball sampling. Participation was voluntary, and all participants provided informed consent prior to the start of interviews. Following the interviews, participants were compensated with Amazon gift cards. All interviews were conducted via video chat. Participants were given the option of either having interviews recorded to create audio transcripts, or having researchers take detailed notes during interviews. We provided the option for note-taking instead of recording as some practitioners expressed concerns about accidentally disclosing confidential company information, although no such disclosures ultimately occurred. In total, we recruited 11 ML practitioners (averaging 5.36 years of experience) across seven big tech companies. Practitioners held titles such as ‘Machine Learning Engineer,’ ‘Machine Learning Researcher,’ or ‘Scientist,’ and were actively involved in building RS. Each participant worked for a company that serves over one billion customers annually. Further details on participants are available in Table 1. This study was supervised by a university Institutional Review Board (IRB).
@@
Table 1. Participant Background Information.:
@@
Interviews ranged between 45 minutes and one hour. We began by asking ice-breaking questions regarding the practitioner’s background, including their years of experience, employer, and the RS that they built or worked with. Practitioners were instructed to be as detailed as possible but to avoid sharing any confidential information such as revenue or upcoming products. Next, we asked questions intended to zero in on practitioner workflows and approaches to fairness. We prepared a set of initial opening questions to encourage practitioners to share their general experiences and provide targets for follow-ups. We then proceeded to a set of follow-up questions about fairness metrics, practitioner time allocation, and cross-team collaboration based on earlier responses. Finally, we invited practitioners to comment on and rate their professional experiences in RS in terms of emotional support, work intensity, and stress level. Our semi-structured interview guide is provided in Appendix Table 3.
@@
We sought to understand how fairness is approached in practice by individuals and teams who develop and deploy large-scale RS. To do so, we analyzed responses on two levels: an affinity diagramming exercise to understand RS practitioner workflows, and a thematic analysis to understand the challenges faced by RS practitioners who seek to improve fairness in the systems that they build, deploy, and maintain.
Affinity Diagramming. First, we conducted an affinity diagramming exercise to map practitioners’ workflows based on interview notes and transcripts, following Yang et al. (2016). We focused on identifying commonalities, and also flagged outlier cases in our notes for further examination. We identified key stages practitioners shared in building RS. We also identified the core set of collaborators and external teams with whom RS practitioners interacted in order to construct a flow model of communications (Wixon et al., 1990). Data were analyzed during collaborative meetings between authors, with interviewers summarizing their sessions and outlining notes and/or transcripts relevant to each session under inspection. We then collaboratively sketched and refined models describing both the general flow of information and data during model development as well as the specific fairness approaches undertaken by practitioners. Findings are described in §4.
Thematic Analysis. Second, we conducted a thematic analysis to identify core challenges faced by practitioners (Braun and Clarke, 2019, 2006). We first converted scripts and notes into items on a Miro web board for aggregation and analysis. Authors conducted two coding passes through all quotes and reflections, developing a large set of initial theme groups. In the second round, authors aggregated thematic groups, merging similar themes and recategorizing elements through negotiation. Lastly, through another open discussion, authors broke down themes that had a significantly larger number of notes into finer granularity and placed any remaining unlabeled items into categories. When finalizing results, authors summarized the high-level themes identified and pulled representative quotes from the grouped items. Findings about technical challenges faced by RS practitioners are described in §5 and findings about organizational challenges arising from RS practitioner interactions with other teams are described in §6.
@@
We are a team of researchers who work on academic and industry teams, and who collectively have expertise in machine learning, human-computer interaction, and algorithmic fairness. Our backgrounds have certainly informed our approach to this study. Our experience on industry teams informed our understanding of the practical constraints that practitioners may face, which influenced the questions we asked in our semi-structured interviews. At the same time, our experience does not extend to working as RS practitioners within technology companies. While this positions us to critically assess RS workflows and challenges as outsiders, we do by definition lack ‘insider’ knowledge about what it is like to work on RS in practice. Finally, we are also all end-users of RS systems. We believe that this dual vantage point as both researchers and users shaped our understanding of the potential real-world impacts of RS, which in turn informed our approach to thematic analysis.
@@ 4. The RS Practitioner Workflow
Figure 1. Recommender System Practitioner Workflow. We visualize RS practitioners’ interactions with the data team managing and monitoring production data access, the legal team auditing and approving specific data access requests, and the fairness (or responsible AI) team providing fairness expertise. Distinct colors—blue for technical (RS practitioners), yellow for legal, red for data, and green for fairness—represent team inputs. Dotted arrows indicate occasional, rather than consistent, participation by a team in certain stages. Offline phases are denoted as gray boxes, while online phases are in purple boxes. Green circles indicate phases in which fairness considerations are commonly incorporated.:
In this section, we describe the RS practitioner workflow. We focus on how RS practitioners integrate fairness work into their labor, both internally (within technical teams) and in collaboration with legal, data, and fairness teams. These findings emerge out of the workflow-centric affinity diagramming exercise described in §3.
@@ 4.1. An Overview of the RS Practitioner Workflow
The RS practitioner workflow is summarized in Figure 1. Practitioners described workflows that evolved from an offline phase (gray boxes in Figure 1) into an online phase (purple boxes in Figure 1) throughout development. First, practitioners coordinate with the legal and data teams to receive access to necessary data and begin the data processing phase. The data team negotiates technical details such as API designs for collecting data and data cleaning, while the legal team reviews data access requests. Practitioners then spend much of their time on model construction and conducting rounds of offline model validation. All agreed that they invested more than half of their work time into these model-centric activities. For example, P5 reported, “This is the core of our daily work. I spent more than 60% time on model work.” Overall, the offline stages of the RS practitioner workflow align with the stages of general ML pipelines, e.g., as described by Suresh and Guttag (2021) and Black et al. (2023), as well as with preferences towards “model work” over “data work” that have been previously documented by Sambasivan et al. (2021).
After offline development, practitioners ship their model and move to an online testing phase, where real-world user behavior data and newly generated user-content interaction data are collected as users engage with recommendations. A fairness team, which all practitioners reported exists in their organizations under various names (such as ‘Responsible AI’), provides feedback to RS practitioners at this stage (see §4.2). Simultaneously, practitioners monitor the model’s performance metrics during the online model service phase, and might revisit the data pre-processing phase to change their training data (e.g., performing feature transformations or asking for more features).
@@ 4.2. How the RS Practitioner Workflow Incorporates Fairness
Fairness assessments and improvements occurred at several points throughout the RS workflow, corresponding to shifts to new development stages (green circles in Figure 1). These transition points were convenient locations to incorporate fairness because they often included interactions between different teams. Within these transition points, the overall process for incorporating fairness considerations into the RS practitioner workflow takes shape. It consists of three major phases, as depicted in Figure 2: the prototyping phase, the internal phase, and the collaborative phase. The prototyping phase co-occurs with the data processing phase of the RS workflow. In the data processing phase, RS practitioners interacted with legal and data teams to obtain access to training data. Concurrently, practitioners started the prototyping phase by hypothesizing about the sources of unfairness that might exist in the training data. Practitioners based hypotheses on their prior experience with RS, and did not generally cite academic research as a source for their hypotheses. Likewise, practitioners reported reusing fairness metrics that they had previously developed to assess training data in this stage. Later, during the offline model validation phase of the RS workflow, practitioners moved to the internal phase of the fairness process, wherein they validated the metrics developed during the prototyping phase and manually searched for unexpected results. Practitioners expressed that there are several rounds of iteration at this stage to find as many instances of potential unfairness as possible prior to model deployment.
Figure 2. Recommender System Practitioner Fairness Process. Technical teams incorporate fairness into their workflows in collaboration with legal, data, and fairness teams. :
@@
After the model is shipped for online testing, the fairness process enters the collaborative phase. This phase is typically the first to include the fairness team, as well as the first phase to include data from real-world users of the RS. In the collaborative phase, the fairness team offers feedback based on their internal guidelines. Most practitioners reported that they had discretion as to whether to integrate feedback from the fairness team. Because practitioners felt that they had limited time and resources available, they sometimes attempted to weigh the relative impact of potential fairness issues against the work required to resolve them. As the owners of an RS within their organizations, practitioners reported numerous ongoing job requirements, such as monitoring the model’s online performance. Practitioners primarily reported being concerned about factors such as uptime, and did not consider fairness to reach a similar level of disruptive emergency necessitating urgent correction. When RS practitioners were able to address fairness issues in deployed models, they often had to integrate user feedback and data to seek out ‘hidden’ types of unfairness (i.e., those that they had not anticipated in the prototyping stage). This process is highly iterative and requires much manual analysis. Practitioners mentioned that, in a typical project, they eventually compose over five new fairness metrics tailored to the specific system and context, even though they typically started with pre-existing metrics from prior projects.
@@ 5. Technical Challenges Arising Within RS Practitioner Teams When Incorporating Fairness
Defining Fairness Defining fairness in specific RS contexts, especially those not subject to pre-existing rules or regulation (11/11) Adapting pre-existing fairness metrics for RS (7/11), and managing the internal documentation involved in those adaptations (8/11)
Considering Multiple Types of Fairness Across Many Groups Managing (conflicting) multi-stakeholder interests (6/11) Prioritizing between different (non-conflicting) types of fairness (8/11) Handling high numbers of user groups (5/11)
Dynamic Fairness Considerations Developing ad hoc metrics to detect unexpected sources of bias, especially when managing real-world user interaction and feedback loops (10/11)
Organizational (Cross-Team) Lack of Time Finding time to work on fairness within technical teams (9/11) and in cross-team collaborations (6/11)
Cross-Team Communication Communicating and collaborating effectively with external teams (7/11)
Table 2. Challenges Faced by RS Practitioners when Incorporating Fairness into their Workflows. Challenges occur within RS teams as well as across teams and at the organization level. The number of practitioners who reported experiencing each type of challenge is shown in parentheses.:
In this section, we describe the key technical challenges faced by RS practitioners who seek to incorporate fairness considerations into their workflows. These findings emerge out of the challenge-centric thematic analysis described in §3. Table 2 details the technical challenges most commonly reported by practitioners.

### v3

@@ 3. Methods
To understand how RS practitioners approach fairness in practice, we conducted 11 semi-structured interviews over a three-week period in 2022. We interviewed RS practitioners from a variety of large companies that deploy highly visible RS. Our interviews were intended to answer two key research questions:
  • RQ1:
    What do the technical and fairness workflows of an RS practitioner look like, and how do they intersect?
  • RQ2:
    What challenges do RS practitioners face when incorporating fairness considerations into their technical workflows?
@@
We recruited participants via social media, direct emails to individuals in our networks or who we had identified based on publicly available information as working in RS, and snowball sampling. Participation was voluntary, and all participants provided informed consent prior to the start of interviews. Following the interviews, participants were compensated with Amazon gift cards. All interviews were conducted via video chat. Participants were given the option of either having interviews recorded to create audio transcripts, or having researchers take detailed notes during interviews. We provided the option for note-taking instead of recording as some practitioners expressed concerns about accidentally disclosing confidential company information, although no such disclosures ultimately occurred. In total, we recruited 11 ML practitioners (averaging 5.36 years of experience) across seven big tech companies. Practitioners held titles such as ‘Machine Learning Engineer,’ ‘Machine Learning Researcher,’ or ‘Scientist,’ and were actively involved in building RS. We specifically focused on technical practitioners because we were interested in understanding how fairness considerations are integrated into the work of building RS and how practitioners handle dual technical and fairness responsibilities. (We discuss the implications of this decision in more detail in §8.1.) Each participant worked for a company that serves over one billion customers annually. Further details on participants are available in Table 1. This study was supervised by a university Institutional Review Board (IRB).
@@
Table 1. Participant Background Information.
A table describing participant background. There are 11 participants (P1-P11). All participants held roles that involved engineering responsibilities, and all but P5 and P7 held roles that involved research responsibilities. Participants had varying years of experience: 3, 4, 4, 5, 6, 5, 4, 5, 7, 3, and 12, respectively.
@@
Interviews ranged between 45 minutes and one hour. Practitioners were instructed to be as detailed as possible but to avoid sharing any confidential information such as revenue or upcoming products. We began by asking ice-breaking questions regarding practitioner background, including their years of experience, employer, and the RS that they built or worked with. Next, we asked questions intended to help us map practitioner workflows (RQ1). We began with open-ended questions (e.g., “could you describe a recent project’s workflow?”, “are there external teams that you interact with regularly?”) and asked more detailed follow-up questions if necessary (e.g., “how do you feel when interacting with the data team?”). We chose to begin with these open-ended, workflow-centric questions in order to develop an understanding of how central fairness work was to practitioners’ broad descriptions of their responsibilities before they were explicitly prompted to consider fairness. Next, we asked about participants’ professional experience with assessing and improving fairness in RS. These included workflow-related questions intended to address RQ1 (e.g., “what amount of time do you spend on fairness in your daily work?”) as well as questions related to potential challenges intended to address RQ2 (e.g., “can you describe a recent time that improving fairness was difficult or failed?”). We also asked about participants’ perceptions of their organizations’ approaches to fairness (e.g., “is fairness in your organization a team effort?”). We asked about these separately to try to disentagle participants’ technical considerations and challenges from organizational ones. Interview quotes included in this paper have been lightly edited for clarity. Our semi-structured interview guide is provided in Appendix Table 3.
@@
We sought to understand how fairness is approached in practice by individuals and teams who develop and deploy large-scale RS. To do so, we analyzed responses on two levels: first, we mapped RS practitioner workflows, and second, we conducted an inductive thematic analysis to understand the challenges faced by RS practitioners who seek to improve fairness in the systems that they build, deploy, and maintain.
RQ1: Mapping Practitioner Workflows. First, we developed a unified mapping of practitioners’ workflows. As our goal was to recognize general patterns among practitioners, we focused on identifying commonalities and flagged outlier cases for discussion and further examination. Data were analyzed during collaborative meetings between authors, with interviewers summarizing their sessions and outlining transcripts and/or notes relevant to each session under inspection. As each transcript or set of notes was discussed, we identified key stages to building an RS that were discussed by practitioners. We also identified the core set of external teams with whom practitioners interacted. We then collaboratively sketched and refined workflow models describing both the general flow of data during model development as well as the specific fairness approaches undertaken by practitioners at each stage of the process. Findings are described in §4.
RQ2: Understanding Practitioner Challenges. Second, we conducted an inductive thematic analysis to identify core challenges faced by practitioners as they sought to incorporate fairness considerations into their technical workflow (Braun and Clarke, 2006). We first converted transcripts and interview notes into items on a Miro web board for aggregation and analysis. Authors conducted two coding passes through all quotes and notes, developing a large set of initial codes related to any mention of challenges that practitioners faced when trying to incorporate fairness into their technical RS workflows. Next, authors collated codes into initial themes. Then, authors reviewed and refined themes, aggregating thematic groups, merging similar themes, and recategorizing elements through negotiation. Lastly, through another open discussion, authors broke down themes that had a significantly larger number of notes into finer granularity and placed any remaining unlabeled items into categories. When finalizing results, authors summarized the high-level themes identified and pulled representative quotes from the grouped items. Findings about technical challenges faced by RS practitioners are described in §5, and findings about organizational challenges arising from RS practitioner interactions with other teams are described in §6.
@@
We are a team of researchers who work on academic and industry teams, and who collectively have expertise in machine learning, human-computer interaction, and algorithmic fairness. Our backgrounds have informed our approach to this study. Our experience on industry teams informed our understanding of the practical constraints that practitioners may face, which influenced the questions we asked in our semi-structured interviews. At the same time, our experience does not extend to working as RS practitioners within technology companies. While this positions us to critically assess RS workflows and challenges as outsiders, we lack ‘insider’ knowledge about what it is like to work on RS in practice. Finally, we are also all end-users of RS systems. This dual vantage point as both researchers and users shaped our understanding of the potential real-world impacts of RS, which in turn informed our approach to thematic analysis.
@@ 4. The RS Practitioner Workflow
Figure 1. Recommender System Practitioner Workflow. We visualize RS practitioners’ interactions with the data team managing and monitoring production data access, the legal team auditing and approving specific data access requests, and the fairness (or responsible AI) team providing fairness expertise. Dotted borders indicate offline phases of the workflow, and solid borders indicate online phases. For each phase of the technical workflow (top row), there are associated phases within the fairness workflow (second row). In each phase, different participants interact (bottom row).
A figure summarizing the RS practitioner technical and fairness workflows. The technical workflow includes the following offline phases: data acquisition and processing, model construction, and model validation; as well as the following online phases: model testing and model service. The fairness workflow co-occurs with the technical workflow. The fairness workflow consists of the prototyping, internal, and collaborative phases. The prototyping phase, which co-occurs with the offline data acquisition and processing and model construction phases, includes the technical team, the legal team, and the data team. In the protoyping phase, the following actions are taken: the technical team makes data requests, the legal team approves or rejects data requests based on regulatory or risk considerations, the data team facilitates data requests via API access, the technical team hypothesizes about sources of unfairness, and the technical team develops or adapts fairness metrics. The internal phase, which co-occurs with the offline model validation phase, includes the technical team. In the internal phase, the technical team validates fairness metrics, conducts iterative fairness testing, and manually searches for previously unknown instances of unfairness. The collaborative phase, which co-occurs with the online model testing and model service phases, includes the technical team, the fairness team, and end users. In the collaborative phase, the fairness team provides feedback based on internal guidelines; end users interact with the RS and provide implicit and explicit feedback; and the technical team manually searches for previously unknown instances of unfairness, determines which fairness and other issues to resolve, modifies fairness metrics, and conducts iterative fairness testing.
In this section, we describe the RS practitioner workflow. We focus on how RS practitioners integrate fairness work internally (within technical teams) and in collaboration with legal, data, and fairness teams.
@@ 4.1. An Overview of the RS Practitioner Workflow
The RS practitioner workflow is summarized in Figure 1. Practitioners described workflows that evolved from an offline into an online phase throughout development.
Offline Development
First, practitioners coordinate with the legal and data teams to receive access to necessary data and begin the data acquisition and processing phase. The data team negotiates technical details such as API designs for data collection as well as data cleaning, while the legal team reviews data access requests. Practitioners then spend much of their time on model construction and conducting rounds of model validation. All participants agreed that they invested more than half of their work time into these model-centric activities. For example, P5 reported, “This is the core of our daily work. I spend more than 60% [of my] time on model work.” Overall, the offline stages of the RS practitioner workflow align with the pre-deployment stages of general ML pipelines as described by Suresh and Guttag (2021) and Black et al. (2023), as well as with preferences towards “model work” over “data work” that have been previously documented by Sambasivan et al. (2021).
Online Development
After offline development, practitioners ship their model and move to a model testing phase, where real-world user behavior data and newly generated user-content interaction data are collected as users engage with recommendations. A fairness team, which all participants reported exists in their organizations under various names (such as ‘responsible AI’), provides feedback to RS practitioners at this stage (see §4.2). Simultaneously, practitioners monitor model performance metrics during the model service phase, and might revisit the data acquisition and pre-processing phase to change their training data (e.g., performing feature transformations or asking for more features). These online stages of the RS practitioner workflow align with the post-deployment stages of general ML pipelines (Suresh and Guttag, 2021; Black et al., 2023). However, while prior work has identified only a small amount of ML fairness research focused on post-deployment efforts (Black et al., 2023), participants described the online phases as a critical—and long-running—part of their workflow. As P2 put it, “We are in [a] loop [where we keep] validating, online testing and improving the model.
@@ 4.2. How the RS Practitioner Workflow Incorporates Fairness
Fairness assessments and improvements occurred at several points throughout the RS workflow, often corresponding to shifts to new development stages. These transition points were convenient locations to incorporate fairness because they often included interactions between different teams. Within these transition points, the overall process for incorporating fairness considerations into the RS practitioner workflow takes shape. It consists of three major phases, as depicted in Figure 1: the prototyping phase, the internal phase, and the collaborative phase.
Prototyping Phase
The prototyping phase co-occurs with the data acquisition and processing phase of the RS workflow. In the data acquisition and processing phase, technical RS practitioners interact with legal and data teams to obtain access to training data. Here, procedural fairness considerations are raised by the first time as the legal team communicates to the technical team which features (e.g., sensitive attributes) should not be used in building RS due to regulatory requirements or company policy. Unlike prior work, which has often found AI practitioners to be at odds with legal teams who see measuring unfairness as a potential legal liability (Costanza-Chock et al., 2022; Wright et al., 2024), we found that participants had overwhelmingly positive opinions of the legal team (for example, P2 saw the legal team as “help[ing] us comply well with our company’s vision and values.”).
Concurrently, practitioners spend time during the prototyping phase hypothesizing about the sources of unfairness that might exist in the training data and developing metrics intended to examine those hypotheses. Participants based hypotheses on their prior experiences with RS, and did not generally cite academic research as a source. P9 stated that even if they were aware of relevant academic research on fairness considerations related to their use case, they still “need to come up with a clear formula and implement it by ourselves—it is really domain-specific.” Instead, participants reported reusing fairness metrics that they had previously developed to assess training data in this stage and emphasized that their initial metrics were “intuition-based” (P3). A similar reliance on bespoke as opposed to standardized fairness metrics and approaches has been noted in prior work on NLP practitioners (Harvey et al., 2025) as well as AI practitioners generally (Madaio et al., 2022, 2024).
Internal Phase
During the offline model validation phase of the RS workflow, practitioners moved to the internal phase of the fairness workflow, wherein they validated metrics developed during the prototyping phase and manually searched for unexpected results. Participants expressed that they undertake several rounds of iteration in this phase to find as many instances of potential unfairness as possible prior to model deployment. In addition, participants emphasized the uncertain nature of the internal phase. According to P1, “It is really hard to predict users’ behaviors—a lot of uncertainties.” P2 described the internal phase bluntly: “We have to look carefully to find those hidden biases. It is not fun.” These comments point to a practical challenge that has not been widely discussed in prior work: because RS are an inherently collaborative technology, testing them requires a model of user behavior. Thus, relying on historical data as a test set, as is commonly done in prediction tasks, is not sufficient in an RS setting.
Collaborative Phase
@@
After the model is shipped for online testing, the fairness workflow enters the collaborative phase. This phase is typically the first to include the fairness team, as well as the first phase to include data from real-world users of the RS. In the collaborative phase, the fairness team offers feedback based on their internal guidelines. Most participants reported that they had discretion as to whether to integrate feedback from the fairness team. Because participants felt that they had limited time and resources available, they sometimes attempted to weigh the relative impact of potential fairness issues against the work required to resolve them:
P1: If [the fairness team’s] feedback is about removing one particular feature, normally we will simply do it. However, if it requires a lot of work, we probably need to have another broader discussion.
As the owners of an RS within their organizations, participants reported numerous ongoing job requirements, such as monitoring online model performance. Participants primarily reported being concerned about factors such as uptime, and did not consider fairness to reach a similar level of disruptive emergency necessitating urgent correction. These types of competing demands on practitioners’ time have been repeatedly noted by prior research (Wang et al., 2024a; Madaio et al., 2022). However, the iterative, manual, and user behavior-dependent aspects of fairness measurement are more specific to RS. When participants were able to address fairness issues in deployed models, they often had to integrate user feedback and data to seek out “hidden” (P2, P7) types of unfairness (i.e., those that they had not anticipated in the prototyping stage). This process is highly iterative and requires much manual analysis. Participants reported that, in a typical project, they eventually composed over five new fairness metrics tailored to the specific system and context, even though they typically started with pre-existing metrics from prior projects.
@@ 5. Technical Challenges Arising Within RS Practitioner Teams When Incorporating Fairness
Defining Fairness Defining fairness in specific RS contexts, especially those not subject to existing rules or regulations Adapting pre-existing fairness metrics for RS, and managing the internal documentation involved in those adaptations
Considering Multiple Types of Fairness Across Many Groups Managing (conflicting) multi-stakeholder interests Prioritizing between different (non-conflicting) types of fairness Handling high numbers of user groups
Dynamic Fairness Considerations Developing ad hoc metrics to detect unexpected sources of bias Managing real-world user interaction and feedback loops
Organizational (Cross-Team) Lack of Time Finding time to work on fairness within technical teams and in cross-team collaborations
Cross-Team Communication Communicating and collaborating effectively with external teams
Table 2. Challenges Faced by RS Practitioners when Incorporating Fairness into their Workflows. Challenges occur within RS teams as well as across teams and at the organization level.
A table summarizing challenges faced by RS practitioners when incorporating fairness into their workflows. Technical challenges (within RS teams) include defining fairness, considering multiple types of fairness across many groups, and dynamic fairness considerations. Defining fairness includes defining fairness in specific RS contexts, especially those not subject to pre-existing rules or regulation, adapting pre-existing fairness metrics for RS, and managing the internal documentation involved in those adaptations. Considering multiple types of fairness across many groups includes managing (conflicting) multi-stakeholder interests, prioritizing between different (non-conflicting) types of fairness, and handling high numbers of user groups. Dynamic fairness considerations include developing ad hoc metrics to detect unexpected sources of bias, especially when managing real-world user interaction and feedback loops. Organizational challenges (cross-team) include lack of time (finding time to work on fairness within technical teams and in cross-team collaborations) and cross-team communication (communicating and collaborating effectively with external teams).
In this section, we describe the key technical challenges faced by RS practitioners who seek to incorporate fairness considerations into their workflows (summarized in Table 2). Overall, we find that defining fairness in the first place is a challenge, echoing findings from prior interview studies conducted on AI and ML practitioners (Madaio et al., 2022). We also find challenges that are more unique to RS, such as that RS practitioners must consider multiple types of fairness across many different stakeholder groups. RS practitioners must not only balance conflicting stakeholder interests, but also prioritize between a large number of different non-conflicting types of fairness. In addition, RS practitioners face unique challenges related to assessing and improving fairness in dynamic settings. Because RS are collaborative systems that are shaped by user behavior, pre-deployment fairness assessments are especially challenging.